OpenStudy (anonymous):

Find the components f and g so that fog=H. H(x)=│Inx-6│ Find the components f and g so that fog=H. H(x)=│Inx-6│ @Mathematics

OpenStudy (anonymous):

Here's one: \[f(x) = |x|\] \[g(x) = \ln(x) - 6\]

OpenStudy (anonymous):

Here's another:\[f(x) = |x - 6|\]\[g(x) = \ln(x)\]

OpenStudy (anonymous):

what is this you wrote?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I've written two functions f and g such that fog(x) = |ln(x) - 6|

OpenStudy (anonymous):

so is that what I would put as my answer? I am stuck on this stuff. lol

OpenStudy (anonymous):

you can put either of the two answers above; both are valid.

OpenStudy (anonymous):

oh ok. Thank you.